OpenStudy (anonymous):

Solve on the interval [0,2π) : 2sin^2 x - 3sin x + 1 = 0

OpenStudy (anonymous):

@amistre64 or anyone else? got any ideas?

OpenStudy (anonymous):

A. x = 2π, x = π/3 B. x = π, x = 2π/3, x = 5π/5 C. x = π/2, x = π/6, x = 5π/6 D. x = 2π, x = π/4, x = 5π/4

jimthompson5910 (jim_thompson5910):

hint: let z = sin(x)

jimthompson5910 (jim_thompson5910):

if you do that, then you can replace every copy of sin(x) with z to get this equation 2z^2 - 3z + 1 = 0

jimthompson5910 (jim_thompson5910):

you would now solve for z, then use those solutions in terms of z to get the solutions in terms of x

OpenStudy (anonymous):

z = 1 z = 1/2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so remember we let z = sin(x), so if z = 1, then sin(x) = 1 ----> x = ???

jimthompson5910 (jim_thompson5910):

if z = 1/2, then sin(x) = 1/2 which means x = ??? or x = ???

jimthompson5910 (jim_thompson5910):

you should get 3 solutions total (on this interval)

OpenStudy (anonymous):

C. x = π/2, x = π/6, x = 5π/6

jimthompson5910 (jim_thompson5910):

very good on all 3

OpenStudy (anonymous):

Can you help with another one then? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

4 csc x + 6 = -2 on the interval [0,2π)

jimthompson5910 (jim_thompson5910):

first isolate csc(x)

jimthompson5910 (jim_thompson5910):

replace it with z if you want like you did last time

OpenStudy (anonymous):

x ≠ 0, z = -2/3

OpenStudy (anonymous):

z = -2/x*

jimthompson5910 (jim_thompson5910):

idk how you got that, if z = csc(x), then z = -2/x doesn't make much sense

OpenStudy (anonymous):

whoops! z = -2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

z = csc(x) as well

jimthompson5910 (jim_thompson5910):

so because of those two facts, we know that csc(x) = -2

jimthompson5910 (jim_thompson5910):

and because csc(x) is really 1/sin(x), we can say 1/sin(x) = -2 sin(x) = -1/2 ... take the reciprocal of both sides

OpenStudy (anonymous):

A. 7π/6, 11π/6 B. π/6, 5π/6 C. π/3, 5π/3 D. 2π/3, 4π/3

jimthompson5910 (jim_thompson5910):

if sin(x) = -1/2, then x = ???? or x = ???

OpenStudy (anonymous):

humm... A?

OpenStudy (anonymous):

B!

jimthompson5910 (jim_thompson5910):

A or B, which one

OpenStudy (anonymous):

A!

jimthompson5910 (jim_thompson5910):

hopefully you're not just randomly guessing

OpenStudy (anonymous):

No B. looked like the first

jimthompson5910 (jim_thompson5910):

i see, they're slightly different

jimthompson5910 (jim_thompson5910):

I recommend you using a unit circle if you're not sure

OpenStudy (anonymous):

B - final answer!

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

if sin(x) = -1/2, then x = 7pi/6 or x = 11pi/6